# RiskCueBench: Benchmarking Anticipatory Reasoning from Early Risk Cues in Video-Language Models

### v1

@@ Abstract
Sha Luo1,*  Yogesh Prabhu2,*  Tim Ossowski1,*  Kaiping Chen1,†  Junjie Hu1,† 1University of Wisconsin–Madison, Madison, WI, USA 2University of California San Diego, San Diego, CA, USA sluo83@wisc.edu, ossowski@wisc.edu, yprabhu2@wisc.edu, kchen67@wisc.edu,junjie.hu@wisc.edu
@@ 1 Introduction
Advanced vision language models (VLMs) have surfaced with remarkable abilities to comprehend complex spatial relationships, temporal sequences, and visual narratives. However, the specific needs of safety critical applications, particularly situa- tional risk assessment and prediction, still differ significantly from those of general video under- standing benchmarks. As shown in Figure 1, most existing benchmarks emphasize post hoc under- standing, where models analyze or describe events after they have fully unfolded, such as answering questions or generating captions given complete visual context caba2015activitynet; Zhou_2025_ICCV; Hong_2025_CVPR.
@@
In contrast, predictive reasoning requires models to anticipate future events from partial and often ambiguous early signals. While current VLMs per- form well on descriptive tasks, their capacity for such anticipatory reasoning remains largely unex- plored. Recent benchmarks targeting next event prediction typically frame evaluation as multiple choice question answering, where models select from predefined outcomes rather than reason freely about future risks wu2024star; wang2025fostering; xun2025rtv; cheng2025video, which lack the open-ended nature of real-world risk. Other work focuses narrowly on traffic accidents via synthetic data or non-reasoning metrics kung2024riskbench; fatima2021global; bao2020uncertainty; hussain2024bi, leaving domains like crowd dynamics unaddressed.
@@ 2 Related Work
Advanced capabilities across temporal visual understanding tasks have been demonstrated by recent developments in video language models. The state-of-the-art in combining language modeling with video encoding mechanisms for intricate temporal reasoning is represented by VideoLLaMA 3 zhang2025videollama, Video LLaVA lin2023video, Qwen-VL bai2025qwen2, and Apollo zohar2025apollo. With the introduction of reinforcement learning techniques for video comprehension by MiMo-RL coreteam2025mimovltechnicalreport, the resolution of long-form video analysis problems by LongVILA-R1 chen2025scaling, and the incorporation of explicit reasoning mechanisms that mimic human cognitive processes by GLM-4.1V-Thinking hong2025glm, specialized reasoning systems have further advanced this field.
@@
Recent work has introduced numerous benchmarks aimed at evaluating the spatio temporal reasoning capabilities of vision language models in video understanding (Table 1). These efforts probe diverse aspects of temporal and spatial cognition, including temporal logic and event ordering, fine grained motion understanding, and dynamic scene interpretation in domains such as egocentric video and autonomous driving. Benchmarks such as TimeLogic swetha2025timelogic and ReXTime NEURIPS2024_32683193 focus on logical and causal reasoning over event sequences, while MotionBench Hong_2025_CVPR and related diagnostic tasks evaluate sensitivity to motion dynamics and temporal direction. MVBench li2024mvbench, VLM4D Zhou_2025_ICCV, and STSBench fruhwirth2025stsbench, assess joint spatial and temporal reasoning across diverse video scenarios, yet many tasks can still be solved using static cues or retrospective access to the full video.
@@ 5 Conclusion
This work was supported by a funded project.

### v2

@@ Abstract
Sha Luo1,*  Yogesh Prabhu2,*  Timothy Ossowski1,*  Kaiping Chen1,†  Junjie Hu1,† 1University of Wisconsin–Madison, Madison, WI, USA 2University of California San Diego, San Diego, CA, USA sluo83@wisc.edu, ossowski@wisc.edu, yprabhu@ucsd.edu, kchen67@wisc.edu, junjie.hu@wisc.edu
@@ 1 Introduction
Advanced vision language models (VLMs) have surfaced with remarkable abilities to comprehend complex spatial relationships, temporal sequences, and visual narratives. However, the specific needs of safety critical applications, particularly situa- tional risk assessment and prediction, still differ significantly from those of general video under- standing benchmarks. As shown in Figure 1, most existing benchmarks emphasize post hoc under- standing, where models analyze or describe events after they have fully unfolded, such as answering questions or generating captions given complete visual context Caba Heilbron et al. (2015); Zhou et al. (2025); Hong et al. (2025a).
@@
In contrast, predictive reasoning requires models to anticipate future events from partial and often ambiguous early signals. While current VLMs per- form well on descriptive tasks, their capacity for such anticipatory reasoning remains largely unex- plored. Recent benchmarks targeting next event prediction typically frame evaluation as multiple choice question answering, where models select from predefined outcomes rather than reason freely about future risks Wu et al. (2024); Wang et al. (2025); Xun et al. (2025); Cheng et al. (2025), which lack the open-ended nature of real-world risk. Other work focuses narrowly on traffic accidents via synthetic data or non-reasoning metrics Kung et al. (2024); Fatima et al. (2021); Bao et al. (2020); Hussain et al. (2024), leaving domains like crowd dynamics unaddressed.
@@ 2 Related Work
Advanced capabilities across temporal visual understanding tasks have been demonstrated by recent developments in video language models. The state-of-the-art in combining language modeling with video encoding mechanisms for intricate temporal reasoning is represented by VideoLLaMA 3 Zhang et al. (2025), Video LLaVA Lin et al. (2023), Qwen-VL Bai et al. (2025), and Apollo Zohar et al. (2025). With the introduction of reinforcement learning techniques for video comprehension by MiMo-RL Team et al. (2025), the resolution of long-form video analysis problems by LongVILA-R1 Chen et al. (2025), and the incorporation of explicit reasoning mechanisms that mimic human cognitive processes by GLM-4.1V-Thinking Hong et al. (2025b), specialized reasoning systems have further advanced this field.
@@
Recent work has introduced numerous benchmarks aimed at evaluating the spatio temporal reasoning capabilities of vision language models in video understanding (Table 1). These efforts probe diverse aspects of temporal and spatial cognition, including temporal logic and event ordering, fine grained motion understanding, and dynamic scene interpretation in domains such as egocentric video and autonomous driving. Benchmarks such as TimeLogic Swetha et al. (2025) and ReXTime Chen et al. (2024) focus on logical and causal reasoning over event sequences, while MotionBench Hong et al. (2025a) and related diagnostic tasks evaluate sensitivity to motion dynamics and temporal direction. MVBench Li et al. (2024), VLM4D Zhou et al. (2025), and STSBench Fruhwirth-Reisinger et al. (2025), assess joint spatial and temporal reasoning across diverse video scenarios, yet many tasks can still be solved using static cues or retrospective access to the full video.
@@ 5 Conclusion
This work was supported by a project funded by American Family Insurance.